# Pre-trained Language Models Learn Remarkably Accurate Representations of Numbers

### v1

@@ Abstract
In response, we propose a novel probing technique that decodes numeric values from input embeddings with near-perfect accuracy across a range of open-source LMs. This proves that after the sole pre-training, LMs represent numbers with remarkable precision. Finally, we find that the embeddings’ preciseness judged by our probe’s accuracy explains a large portion of LM’s errors in elementary arithmetic, and show that aligning the embeddings with the pattern discovered by our probe can mitigate these errors.
@@
Marek Kadlčík   Michal Štefánik   Timothee Mickus   Michal Spiegel   Josef Kuchař TransformersClub @ Faculty of Informatics, Masaryk University Language Technology, University of Helsinki Kempelen Institute of Intelligent Technologies
@@ 2 Related Work
Another line of work investigates how models learn to represent and process numerical information. Nanda et al. (2023) show that a transformer with one-hot encoding trained from scratch on modular addition discovers Fourier basis and its computation is interpretable in trigonometric functions. Kantamneni and Tegmark (2025) discover an analogous circuitry for (non-modular) addition in a general pretrained language model, and find that its intermediate representations combine both linear and periodic components, reminiscent of a helix structure. Zhou et al. (2024) further identifies subcomponents of the addition circuitry implemented by the attention mechanism and feedforward layers. Zhu et al. (2025) demonstrate that hidden states of pretrained language models can be approximately decoded with a linear (or multi-layer) probe to estimate the logarithm of the number value. Although the probe outputs correlate with the target value, decoding achieves low accuracy.
@@ 3 Recovering numerical information from number embeddings
We evaluate the probes in Equations 1, 2, 3 and 4 using a cross-validation setup with 20 folds. We report their accuracy measured by rounding the output of the regression probes Equations 1 and 2 to the nearest integer, or by retrieving the index of the row in $\mathbf{S}$ or $\mathbf{B}$ that maximizes the output distribution of the classifier probes Equations 3 and 4. We control the validity of our probes by ensuring that they reach an accuracy of 0 for standard Gaussian vectors as well as for a random permutation of the embeddings. Parameters for regressions are estimated using a least-squares algorithm; whereas our classifiers’ parameters are optimized with Adam with a learning rate of $0.0001$, weight decay of $0.001$, and $\beta=(0.9,0.999)$. We choose a hidden dimension of 100. The classifiers are optimized to distinguish output only between training tokens, and during testing, must choose between all tokens. The probes are optimized until loss converges on a validation split separate from the testing split.
@@ Results.
We summarize performances, measured in terms of accuracy, in Figure 1. Crucially, we are almost systematically able to retrieve the integer value corresponding to the embedding’s number with very high accuracy. Another salient observation is that $f_{\sin}$ consistently outperforms all other probe architectures including the regression probe used in previous work of Zhu et al. (2025), contradicting their finding that LMs learn to encode numbers linearly. Explaining the success of the Fourier basis, we note that other prior literature has suggested that sinusoidal features are used for arithmetic computation in LMs (Zhou et al., 2024). Adding onto this, we can also stress that, qualitatively, most of the models’ whose number embeddings we survey here exhibit wave-like patterns in a PCA projection and have sparse Fourier transform, confirming regularity in the hidden structure. See Figures 3 and 4 in Section A.1 for visualizations of PCA and its Fourier transform. Notably, OLMo 2 32B is the only model with low resemblance of the pattern, which is consistent with the low performance of its sinusoidal probe.
@@ 4 Leveraging numerical information from number embeddings
Motivations.
Having established that number embeddings do encode retrieval numerical information about the integers they represent, we now turn to how this numerical information is leveraged by LMs to perform arithmetic tasks. We study the zero-shot performances of a subset of our models on addition and subtraction tasks. We define our addition task as taking any pair of integers $x_{1}$, $x_{2}$ such that $5000<x_{i}<$ 500 as input, and computing the expected output $x_{1}+x_{2}$. The subtraction task is defined by taking as inputs any pair $x_{1}$, $x_{2}$ such that $10000<x_{2}<x_{1}<$ 1000, and computing the expected output $x_{1}-x_{2}$.
Performances.
To perform the arithmetic tasks, we conduct minimal prompt engineering: we systematically evaluate a handful of natural language prompts for their accuracy in a zero-shot setting on the 1B models, and then select the highest-performing for subsequent analyses. All prompts are listed in Appendix B, see Figure 6(a) for addition and Figure 6(b) for subtraction.
OLMo2 1B
OLMo2 7B
OLMo2 13B
OLMo2 32B
Llama 3 1B
Llama 3 3B
Llama 3 8B
Phi 4 15B
Add. $22.21$ $1.12$ $0.21$ $0.05$ $2.58$ $0.45$ $0.24$ $0.0$
Sub. $28.08$ $0.36$ $0.17$ $0.37$ $1.42$ $0.15$ $0.01$ $0.0$
An overview of the error rates from the LMs we study is listed in Table 1. As is apparent, most models achieve high degrees of performance (except for OLMO 2 1B); we also observe a trend towards fewer errors for models with more parameters.
(a) Llama 3 1B, addition.
(b) Llama 3 1B, subtraction.
(c) OLMo 2 1B, addition.
(d) OLMo 2 1B, subtraction.
Model behavior.
To better explain the behavior of the LMs, we conduct a simple circuit analysis and a feature attribution experiment using integrated gradients (Sundararajan et al., 2017). For convenience, we focus on the two smaller models in our panel. OLMo 2 1B and Llama 3 1B. Both experiments suggest one major difference between operand pairs leading to failure and to success: the probability assigned by the LLM to the predicted output token tends to be statistically lower when the model produces an incorrect output, see Figure 2. We also observe the same subset of heads being activated for failure and success on the arithmetic task; further details are available in Section A.2. Besides the usefulness of this difference in probability mass for diagnostic purposes, these experiments also suggest a difference in degree rather than kind between failures and successes.
@@
Add. 39.120 337 030 102 25639.12033703010225639.120\,337\,030\,102\,25639.120 337 030 102 256 21.335 234 022 333 09721.33523402233309721.335\,234\,022\,333\,09721.335 234 022 333 097 41.221 374 045 801 52541.22137404580152541.221\,374\,045\,801\,52541.221 374 045 801 525 4.607 508 532 423 2084.6075085324232084.607\,508\,532\,423\,2084.607 508 532 423 208 1.243 339 253 996 447 61.24333925399644761.243\,339\,253\,996\,447\,61.243 339 253 996 447 6 8.006 535 947 712 4188.0065359477124188.006\,535\,947\,712\,4188.006 535 947 712 418
Sub. $9.0625$ 29.509 100 937 672 3729.5091009376723729.509\,100\,937\,672\,3729.509 100 937 672 37 26.832 018 038 331 45526.83201803833145526.832\,018\,038\,331\,45526.832 018 038 331 455 0.393 700 787 401 574 770.393700787401574770.393\,700\,787\,401\,574\,770.393 700 787 401 574 77 1.041 666 666 666 666 51.04166666666666651.041\,666\,666\,666\,666\,51.041 666 666 666 666 5 $0.0$
Table 2: Proportion of errors on arithmetic tasks involving misrepresented numbers (%).
@@
To assess how numerical information and arithmetic reasoning are linked, we evaluate whether the errors we see in these arithmetic tasks can be linked to defects of the number embeddings used as inputs. This entails verifying for every pair of operands $x_{1}$, $x_{2}$ that leads to failure whether either of $x_{1}$ or $x_{2}$ also leads to a failure in the most accurate probes set up in Section 3, viz. the $f_{\sin}$ probes. The proportion of errors that involve operands not well captured by the $f_{\sin}$ probe varies greatly, depending on the model and task; ranging from more than 40% of all erroneous pairs to 0%; percentages tend to be higher for models that the $f_{\sin}$ probe did not capture well. This suggests that the quality of representations is a factor impacting arithmetic capabilities in models.
@@
We hypothesize that embeddings of tokens that our probes can not correctly decode might diverge from the model’s robust representation scheme and thus, contribute to errors in arithmetic tasks. With this motivation, we test whether a direct intervention on the embeddings of these tokens can improve models’ performance on arithmetic. In practice, we start from the $f_{\sin}$ probes described in Equation 3 and trained for Llama 3 1B and freeze all probe parameters. We then perform gradient descent to optimize the embeddings of all incorrectly decoded tokens (namely 0, 4, 977 and 999) with respect to the probe decoding loss, aiming to align those tokens with the overall pattern discovered by the probe. We finally measure how this embedding intervention impacts model accuracy on addition and multiplication tasks involving these four tokens as one of the inputs or expected outputs, using the model’s best-performing template (a set of our experimental templates is listed in Table 6(a)).
We find that in additions involving these assumably divergent tokens, our intervention reduces 26% of errors (from 17.6% to 13.0%). In multiplications, our intervention brings error reduction by 9.4% (from 8.5% to 7.7%). This experiment, while of an anecdotal scale determined by a low error rate of our probes, shows that more accurate probes of models’ representations can also guide direct refinements of models’ possibly imprecise embeddings, aligning them with the model’s general hidden structure and bringing observable improvements in accuracy of model’s predictions.
@@ 5 Conclusion
In this paper, we have inspected the embedding representations for number tokens across a range of widely used open-source LMs. Our observations consolidate a growing body of studies showcasing how LMs learn sinusoidal hidden structure in number representations. Building upon this observation, we design a probing method leveraging this structure that decodes LMs’ embeddings with near-perfect accuracy across multiple models, demonstrating that the quality of numeric representations in pretrained LMs was strongly underestimated in previous work. Still, we find a model (OLMo 2 32B) that deviates from this pattern, calling into question the generalizability of the conclusions of works such as Zhou et al.’s (2024). Finally, we show that the preciseness of embeddings relative to the sinusoidal pattern can explain a proportion of practical errors on arithmetic tasks. Furthermore, we demonstrate improved accuracy on those tasks by aligning imprecise embeddings to the model’s learned embedding pattern.
@@ Acknowledgements
[…]
@@
First, our probing method, though highly effective for many models, relies on an assumed hidden structure of models’ learned representations, and therefore expects a priori knowledge of models’ mechanics. This necessarily limits the applicability of our approach to models where a known structure exists, e.g., it is not applicable for OLMo 2 32B.
@@ A.1 Wave-like patterns in embeddings
Figure 3 displays the sinusoidal patterns in Llama 3 70B and OLMo2 13B after re-alignment and de-trending with PCA. For clarity, we only include the first 16 principal components.
@@
(b) Llama 3 70B
(c) OLMo 2 1B
(d) OLMo 2 32B
@@
In Figure 5, we present an overview of head-level attribution of the logits in Llama 2 1B. The same heads in Layers 13 through 15 appear activated in all cases, playing the same inhibitor and booster roles. Incorrectly performed addition leads to a noisier overall pattern. Remarkably, we observe that activity occurs in the latter stages of the model, whereas input embeddings (layer 0) already contain precise numeric information, as per our probing experiments. This delayed processing may explain some of the errors we observe, despite the high accuracy of our probes in Section 3.
@@ Appendix B Experimental details
$x_{1}$+$x_{2}$ equals to
The result of $x_{1}$+$x_{2}$ is
The result of $x_{1}$ plus $x_{2}$ is
The result of $x_{1}$ plus $x_{2}$ =
The result of $x_{1}$ plus $x_{2}$ =
$x_{1}$ plus $x_{2}$ equals to
$x_{1}$+$x_{2}$=
$x_{1}$ plus $x_{2}$ equals
$x_{1}$ plus $x_{2}$ is equal to
$x_{1}$+$x_{2}$ equals
$x_{1}$+$x_{2}$ is equal to
$x_{1}$ plus $x_{2}$ equals
$x_{1}$ plus $x_{2}$ is equal to

### v2

@@ Abstract
In response, we propose a novel probing technique that decodes numeric values from input embeddings with near-perfect accuracy across a range of open-source LMs. This proves that after the sole pre-training, LMs represent numbers with remarkable precision. Finally, we find that the embeddings’ precision, judged by our probe’s accuracy, explains a large portion of LM’s errors in elementary arithmetic, and show that aligning the embeddings with the pattern our probes discover can mitigate these errors.
@@
Marek Kadlčík♣† Michal Štefánik♡♣∗† Timothee Mickus♡† Michal Spiegel♣♢ Josef Kuchař TransformersClub @ Faculty of Informatics, Masaryk University Language Technology, University of Helsinki Kempelen Institute of Intelligent Technologies Core contributors
**footnotetext: From October 2025 at National Institute of Informatics Research and Development Center for Large Language Models, Tokyo, Japan
@@ 2 Related Work
Another line of work investigates how models learn to represent and process numerical information. Nanda et al. (2023) show that a transformer with one-hot encoding trained from scratch on modular addition discovers Fourier basis and its computation is interpretable in trigonometric functions. Kantamneni and Tegmark (2025) discover an analogous circuitry for (non-modular) addition in a general pretrained language model, and find that its intermediate representations combine both linear and periodic components, reminiscent of a helix structure. Zhou et al. (2024) further identifies subcomponents of the addition circuitry implemented by the attention mechanism and feedforward layers. Zhu et al. (2025) demonstrate that hidden states of pretrained language models can be approximately decoded with a linear (or multi-layer) probe to estimate the logarithm of the number value. Although the probe outputs correlate with the target value, decoding achieves low accuracy. Recently, Levy and Geva (2025) show success in recovering the values of digits from internal representations of intermediate layers, hinting on a more generalized, circular pattern in representations of numbers.
@@ 3 Recovering numerical information from number embeddings
We evaluate the probes in Equations˜1, 2, 3 and 4 using a cross-validation setup with 20 folds. We report their accuracy measured by rounding the output of the regression probes Equations˜1 and 2 to the nearest integer, or by retrieving the index of the row in $\mathbf{S}$ or $\mathbf{B}$ that maximizes the output distribution of the classifier probes Equations˜3 and 4. We control the validity of our probes by ensuring that they reach an accuracy of 0 for standard Gaussian vectors as well as for a random permutation of the embeddings. Parameters for regressions are estimated using a least-squares algorithm; whereas our classifiers’ parameters are optimized with Adam with a learning rate of $0.0001$, weight decay of $0.001$, and $\beta=(0.9,0.999)$. We choose a hidden dimension of 100. The classifiers are optimized to distinguish output only between training tokens, and during testing, must choose between all tokens. The probes are optimized until loss converges on a validation split separate from the testing split.
We release an implementation and training recipes for the new probes, including all configurations we use, in the project’s GitHub repository. [^1] https://github.com/prompteus/numllama
@@ Results.
We summarize performances, measured in terms of accuracy, in Figure˜1. Crucially, we are almost systematically able to retrieve the integer value corresponding to the embedding’s number with very high accuracy. Another salient observation is that $f_{\sin}$ consistently outperforms all other probe architectures including the regression probe used in previous work of Zhu et al. (2025), contradicting their finding that LMs learn to encode numbers linearly. Explaining the success of the Fourier basis, we note that other prior literature has suggested that sinusoidal features are used for arithmetic computation in LMs (Zhou et al., 2024). Adding onto this, we can also stress that, qualitatively, most of the models’ whose number embeddings we survey here exhibit wave-like patterns in a PCA projection and have sparse Fourier transform, confirming regularity in the hidden structure. See Figures˜2 and 3 in Section˜A.1 for visualizations of PCA and its Fourier transform. Notably, OLMo 2 32B is the only model with low resemblance of the pattern, which is consistent with the low performance of its sinusoidal probe.
Analysis.
To verify that our sin-base probes indeed reach their superior accuracy by learning to extract a generalized, sin-like representation from models’ representations, we analyse the encoded representations that trained sin probes produce as the output of $\mathbf{W}_{\mathrm{in}}$. We experiment with two training settings: (i) using L1 regularization — encouraging sparsity, and (ii) using L2 regularization — encouraging the employment of a broader scale of input features. We note that in both of these settings, the probes achieve almost identical generalization capacity as assessed by their accuracy on unseen inputs (embeddings of numbers).
Figure 6 in Appendix A.3 displays the resulting representations for model embeddings of Llama3 1B associated with different numeric values. We can observe that the L1- and L2-regularized probes learn a substantially distinct representational pattern. We hypothesize that a main difference between probes trained with different regularizations is that the L1 probe learns to follow a broader scale of distinct frequencies, while the L2 probe follows similar frequencies shifted by a different constant. Nevertheless, in both of the cases, the probe learns a projection into a wave-like pattern across input numbers, thus successfully following their injected inductive bias.
@@ 4 Leveraging numerical information from number embeddings
Motivation.
Having established that number embeddings do encode retrieval numerical information about the integers they represent, we now turn to how this numerical information is leveraged by LMs to perform arithmetic tasks. We study the zero-shot performances of a subset of our models on addition and subtraction tasks. We define our addition task as taking any pair of integers $x_{1}$, $x_{2}$ such that $0<x_{i}<500$ as input, and computing the expected output $x_{1}+x_{2}$. The subtraction task is defined by taking as inputs any pair $x_{1}$, $x_{2}$ such that $0<x_{2}<x_{1}<1000$, and computing the expected output $x_{1}-x_{2}$.
Performance.
To perform the arithmetic tasks, we conduct minimal prompt engineering: we systematically evaluate a handful of natural language prompts for their accuracy in a zero-shot setting, and then select the highest-performing for subsequent analyses. Due to computational costs, we ignore the two largest models (OLMo2 32B and Llama 3 70B). All prompts are listed in Appendix˜B, see Figure˜7(a) for addition and Figure˜7(b) for subtraction.
@@
Phi 4 15B
Add. $21.39$ $1.12$ $0.17$ $2.58$ $0.45$ $0.25$ $0.0$
Sub. $28.12$ $0.36$ $0.16$ $1.43$ $0.03$ $0.01$ $0.0$
An overview of the error rates from the LMs we study is listed in Table˜1. As is apparent, most models achieve high degrees of performance (except for OLMO 2 1B); we also observe a trend towards fewer errors for models with more parameters.
@@
To assess how numerical information and arithmetic performance are linked, we evaluate whether the errors we see in these arithmetic tasks are associated with defects of the number embeddings used as inputs.
We measure the error rate on the downstream addition and subtraction task in two separate cases – in the first case, both input tokens are decodable by the probe, in the second case, at least one value is not. The results can be seen in Table˜2.
OLMo2 1B
OLMo2 7B
OLMo2 13B
Llama 3 1B
Llama 3 3B
Llama 3 8B
Addition
  decodable $20.30$ $0.98$ $0.16$ $2.48$ $0.46$ $0.24$
  undecodable $23.33$ $1.84$ $0.20$ $14.86$ $0.28$ $0.28$
Subtraction
  decodable $24.61$ $0.32$ $0.19$ $1.43$ $0.03$ $0.01$
  undecodable $31.45$ $0.53$ $0.13$ $0.90$ $0.04$ $0.0$
Table 2: Downstream arithmetic error rate (%, $\downarrow$) given that (1) all tokens are decodable, and (2) at least one token is non-decodable. Results are measured on all possible input combinations. Phi is omitted because it does not make errors.
The results show that models tend to make more errors when the input embeddings are misaligned with the pattern used by the probe, as undecodable inputs lead to higher error rates in 8 out of 12 configurations. The effect is more prominent for models with substantial error rates, such as OLMo2 1B.
@@
We hypothesize that embeddings of tokens that our probes can not correctly decode diverge from the model’s robust representation scheme and thus contribute to errors in arithmetic tasks. With this motivation, we test whether a direct intervention on the embeddings of these tokens can improve models’ performance on arithmetic. In practice, we start from the $f_{\sin}$ probes described in Equation˜3 and trained for Llama 3 1B and freeze all probe parameters. We then perform gradient descent to optimize the embeddings of all incorrectly decoded tokens (namely 0, 4, 977 and 999) with respect to the probe decoding loss, aiming to align those tokens with the overall pattern discovered by the probe.
We finally measure how this embedding intervention impacts model error rate on addition and multiplication tasks involving these four tokens as one of the inputs or expected outputs, using the model’s best-performing template (a set of our experimental templates is listed in Table 7(a)).
We find that in additions involving these assumably divergent tokens, our intervention reduces 26% of errors (from 17.6% to 13.0%). In multiplications, our intervention brings error reduction by 9.4% (from 8.5% to 7.7%). This experiment, while of an anecdotal scale determined by a low error rate of our probes, shows that more accurate probes of models’ representations can also guide direct refinements of models’ possibly imprecise embeddings, aligning them with the model’s general hidden structure and bringing improvements in accuracy of the model’s predictions.
@@ 5 Conclusion
In this paper, we have inspected the embedding representations for number tokens across a range of widely used open-source LMs. Our observations consolidate a growing body of studies showcasing how LMs learn sinusoidal hidden structure in number representations. Building upon this observation, we design a probing method leveraging this structure that decodes LMs’ embeddings with near-perfect accuracy across multiple models, demonstrating that the quality of numeric representations in pretrained LMs was strongly underestimated in previous work. Still, we find a model (OLMo 2 32B) that deviates from this pattern, calling into question the generalizability of the conclusions of works such as Zhou et al.’s (2024). Finally, we show that the preciseness of embeddings relative to the sinusoidal pattern can explain a proportion of practical errors on arithmetic tasks, especially when models fail to align closely with this sinusoidal pattern.
Furthermore, we demonstrate improved accuracy on those tasks by aligning imprecise embeddings to the model’s learned embedding pattern. To some extent, our findings curtail the validity of offloading approaches for numerical reasoning (Schick et al., 2023; Kadlčík et al., 2023): showing that their initial premise — of models not learn accurately representations of numbers — is incorrect. We hope that our findings will motivate future work to rigorously compare relative advantages of tool-using models in terms of computational efficiency, and challenge future work towards the data Štefánik et al. (2024) and architecture refinements Spiegel et al. (2025) accelerating more efficient learning of accurate representations of exact elements of language.
@@ Acknowledgements
This work is supported by the Research Council of Finland through project No. 353164 “Green NLP – controlling the carbon footprint in sustainable language technology”.
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
$\bigstar$
This project has received funding from the European Union’s (EU) Horizon Europe research and innovation programme under Grant agreement No. 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee (grant number 10052546). The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the EU.
@@
First, our probing method, though highly effective for many models, relies on an assumed hidden structure of models’ learned representations, and therefore expects a broad a priori understanding of models’ representation space. This necessarily limits the applicability of our approach to models where a known structure exists; Our results aim to show that some language models indeed do exhibit alternativel encoding schemes, exemplified by OLMo 2 32B that, ableit being highly accurate in arithmetics, can not be accurately probed by our sinusoidal probes.
@@
Fourth, our analysis targets model embeddings. It is thus limited to single-token representations, and does not address the inner mechanisms of numeric information processing in large language model. This area also calls for further research.
@@ A.1 Wave-like patterns in embeddings
Figure˜2 displays the sinusoidal patterns in Llama 3 70B and OLMo2 13B after PCA dimensionality reduction. For clarity, we only include the first 16 principal components.
@@
(b) Llama 3 3B
(c) Llama 3 8B
(d) Llama 3 70B
(e) OLMo 2 1B
(f) OLMo 2 7B
(g) OLMo 2 13B
(h) OLMo 2 32B
(i) Phi 4 15B
@@
Model behavior.
To better explain the behavior of the LMs, we conduct a simple circuit analysis and a feature attribution experiment using integrated gradients (Sundararajan et al., 2017). For convenience, we focus on the two smaller models in our panel. OLMo 2 1B and Llama 3 1B.
Both experiments suggest one major difference between operand pairs leading to failure and to success: the probability assigned by the LLM to the predicted output token tends to be statistically lower when the model produces an incorrect output, see Figure˜4. We also observe the same subset of heads being activated for failure and success on the arithmetic task. Besides the usefulness of this difference in probability mass for diagnostic purposes, these experiments also suggest a difference in degree rather than kind between failures and successes.
(a) Llama 3 1B, addition.
(b) Llama 3 1B, subtraction.
(c) OLMo 2 1B, addition.
(d) OLMo 2 1B, subtraction.
@@
In Figure˜5, we present an overview of head-level attribution of the logits in Llama 2 1B. The same heads in Layers 13 through 15 appear activated in all cases, playing the same inhibitor and booster roles. Incorrectly performed addition leads to a noisier overall pattern. Remarkably, we observe that activity occurs in the latter stages of the model, whereas input embeddings (layer 0) already contain precise numeric information, as per our probing experiments. This delayed processing may explain some of the errors we observe, despite the high accuracy of our probes in Section˜3.
A.3 Analysis of sin-base probes’ learned representations
In Figure 6, we can see that our newly proposed sin-like probes indeed learn to project input embeddings of models into an expected, generalized wave-like representation.
(a) L1 regularization
(b) L2 regularization
Figure 6: Hidden representations of sin-base probes for numeric input embeddings of Llama 3 1B model, after training with different regularization strategies show that our sin-base probes learn to project numeric embeddings into a generalized, wave-like representation used as target inductive bias.
@@ Appendix B Experimental details
1 $x_{1}$$x_{2}$ equals to
2 The result of $x_{1}$$x_{2}$ is
3 The result of $x_{1}$ plus $x_{2}$ is
4 The result of $x_{1}$ plus $x_{2}$ =
5 The result of $x_{1}$ plus $x_{2}$ =
6 $x_{1}$ plus $x_{2}$ equals to
7 $x_{1}$$x_{2}$=
8 $x_{1}$ plus $x_{2}$ equals
9 $x_{1}$ plus $x_{2}$ is equal to
10 $x_{1}$$x_{2}$ equals
11 $x_{1}$$x_{2}$ is equal to
12 $x_{1}$ plus $x_{2}$ equals
13 $x_{1}$ plus $x_{2}$ is equal to